Mir Status Report 11 April 1997

Mir 23 / NASA 4 Status (edited)
Date: 11 April 1997 / Mission Day: 90
Mir Station System/Subsystem Activities
IFM's (entire crew)

Mission Science Payload Activities
Sleep Study (NASA 4)
CSST (NASA 4)

PAO: Press Conference held today between Jerry and reporters from JSC/KSC. Comm quality was spotty with a late start due to Mir-related discussions and a two-minute comm dropout. It was also terminated two minutes early by the SFD for further systems discussion. The MOST team will discuss the abruptness of the cutoff with Mr. Blagov.

Mir Systems
Per a conversation with V. Blagov, the VGK in Kvant 1 is still leaking, and therefore, will not provide cooling to the Vosdukh. Mr. Blagov said the plans are to operate the Vosdukh without cooling. He also said the Vosdukh system used in the design and ground testing of the BURAN had no thermal interface, so there is no concern with operating the Vosdukh for at least 30 days.

In this mode of operation, the system will dump moisture out of the cabin, so

the 30 day limit is related to the lower limit of ppH2O. At 09:49 DMT the ppH20 was 13.92 mm Hg. The Vosdukh was turned on in this mode before the 11:16 DMT pass, at which the crew reported nominal operations, i.e. no alarms. The nominal temperature limit for the CO2 absorbing beds, without degradation of efficiency, is 23C. If these beds are exposed to 50C air, the system will stop working (the CO2 removal beds will not remove CO2). The plans are to operate the Vosdukh in Kvant 1 in this mode until the leak in VGK can be found and repaired. The crew will be working on this activity until the leak can be found.

The crew is systematically searching for the leak in VGK. This includes isolating segments of lines and monitoring for a pressure drop. The crew was also asked to carefully inspect the hydroconnectors.

The Antares transmitter (which will be ready for shipment to the US on 20th April) will replace one of the two transmitters. One is hard failed, and the other is operational, but because of the lack of cooling (from KOB-1) to it, it will only operate for 10 minutes.

The Elektron system in the Kvant-2 was started last night, but only remained on for a few minutes each time. Mr. Blagov said that the ground thinks that the problem is a faulty fluid sensor in the control unit. Plans were to replace this sensor with an identical unit from the Elektron in the Kvant-1. Meanwhile, fluid pressure in the Elektron appeared to hold, and the crew was instructed to continue to attempt to operate the Elektron, cycling power back if it continues to kick back off "as many times as it takes."

Late today, the crew complained they had too much to do, so the ground instructed them to suspend repair work on the Elektron for now, and focus on finding the leak in the VGK.

The crew also reported that the Urine Processing System had shut down, with a malfunction light showing "change separator." This had occurred last night, apparently due to failure of the MNR fan separator. This was replaced last night, but the system shut down again today. The ground has suggested a sequence of steps to manually override the automatic shutdown system, to be followed each time the crew uses the Urine Processing System.

The timelines for Saturday and Sunday have time allocated to the entire crew for maintenance work.

Environmental Parameters (Information from 10 April)

Parameter Min. Max. Nominal Range

Pressure (mmHg) Base Block 748 755 660 - 850
ppO2 (mmHg) 156 159 140 - 200
ppCO2 (mmHg) 4.3 4.7 up to 8.0
ppH20 (mmHg) 14.8 15.9 8 - 14
Humidity (percentage) 37 40 30 - 75
Temperature (C) base block (BB) 33.8 34.5 18 - 28
Temp (C) docking/transfer module 22.2 22.6 18 - 28
Temp (C) work area Krystall 25.2 25.5 18 - 28
Temp (C) sleep area Krystall 26.2 27.8 18 - 28
Temp (C) Spektr 26.0 26.1 18 - 28
Temp (C) Kvant II 20.7 21.5 18 - 28

Planning

MIR 23/NASA 4 MD 62/91
Condensate Presence Check of Mir (Mir 23)
Wet Cleanup (entire crew)
A/G Audio with Friends in celebration of Cosmonauts Day (entire crew)
Rest (entire crew)
Audio PFC (Mir 23)
Sleep Study - Final Ops and Stowage (NASA 4)
BTS Filter Check (NASA 4)

MIR 23/NASA 4 MD 63/92
KOB-2 Loop Leak Repair (Mir 23)
Audio PFC (NASA 4)
Rest (entire crew)

MIR 23/NASA 4 MD 64/93
KOB-2 Loop Leak Repair (Mir 23)
QUELD Sample E-13 (NASA 4)
DOZA - TLD Processing (NASA 4)

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on April 11, 1997 12:32 PM.

Mir Status Report 10 April 1997 was the previous entry in this blog.

Mir Status Report 12 April 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.